A tractor of mass 5.0x10^3kg is used to tow a car of 2.5x10^3kg. the tractor moved with a speed of 3.0m/s just before the towing rope became taut.calculate the speed of the tractor immediately the rope becomes taut.
1
Expert's answer
2020-03-02T10:21:13-0500

From the conservation of momentum:


"Mv=(M+m)u"

"(5000)3=(5000+2500)u"

"u=2.0\frac{m}{s}"